# Guided Decoding for Robot Motion Generation and Adaption

### v1

@@ Abstract
We address motion generation for high-DoF robot arms in complex settings with obstacles, via points, etc. A significant advancement in this domain is achieved by integrating Learning from Demonstration (LfD) into the motion generation process. This integration facilitates rapid adaptation to new tasks and optimizes the utilization of accumulated expertise by allowing robots to learn and generalize from demonstrated trajectories.
We train a transformer architecture on a large dataset of simulated trajectories. This architecture, based on a conditional variational autoencoder transformer, learns essential motion generation skills and adapts these to meet auxiliary tasks and constraints. Our auto-regressive approach enables real-time integration of feedback from the physical system, enhancing the adaptability and efficiency of motion generation. We show that our model can generate motion from initial and target points, but also that it can adapt trajectories in navigating complex tasks, including obstacle avoidance, via points, and meeting velocity and acceleration constraints, across platforms.
@@ I Introduction
Motion generation is a challenging topic for robot arms with high degrees of freedom (DoF). One main challenge is to generate smooth motion trajectories to solve a given task while ensuring the robot and its environment’s safety, and respecting the robot’s embodiment constraints. This problem is especially complex when the task involves auxiliary objectives beyond reaching a target point, such as avoiding obstacles, passing via specified points, or imposing velocity or acceleration constraints. Another particularly intriguing aspect of motion generation is the incorporation of learning from demonstration (LfD) within the same process. Allowing users or agents to present a guiding trajectory based on prior knowledge helps the rapid development of robotic applications. This capability is crucial for adapting to novel tasks and leveraging accumulated expertise effectively. Motion primitives offer a promising avenue for achieving this goal [1, 2, 3, 4]. By encapsulating fundamental motion behaviors into primitives, the robot can learn and generalize from demonstrations, promoting versatility and adaptability in complex tasks. Integrating learning from demonstration with the challenges of motion generation presents an exciting direction for advancing the capabilities of robot arms with high degrees of freedom.
@@
In this work, we propose a framework for learning motion generation. We create a large dataset of robot motion trajectories with multiple simulated robots and train a transformer architecture to learn to generate motion based on given initial and target points. We design a Transformer conditional variational autoencoder. The trained generative model encapsulates basic motion generation skills based on the training data. We further propose an approach to efficiently adapt the generated motion to satisfy auxiliary tasks and constraints such as obstacle avoidance, via points and velocity constraints. The whole process is auto-regressive allowing for integrating feedback from the real system during motion generation. Our approach allows for generating motion based on an initial and target point only, or to imitate a full trajectory given by a user or external agent. Our experiments show the usability of our method for motion generation, under via point constraints, obstacles, and velocity and acceleration constraints. We also show the benefits of training with data collected using various robots. Additionally, we show the ability to imitate and adapt user-given trajectories to a given task.
@@ II Methods
In this section, we present a Transformer-based autoencoder model in Sec. II-A to II-C, followed by an introduction to adaptation methods based on an autoregressive decoder in Section II-D and II-E (see Fig. 1).
@@ II-A Inference model for the infilling task
where the latent variables $z$ are expected to encode latent features of paths with starting point $x_{1}$ and endpoint $x_{T}$. We learn this by minimizing the negative ELBO
@@
Here $p_{\theta}(z|x_{1,T})$ is the conditional prior model, $p_{\theta}(x_{2:T-2}|x_{1,T},z)$ is the decoder/generation model, and $q_{\phi}(z|\operatorname{x}_{1:T})$ is the encoder/recognition model. To generate a trajectory, we sample the latent of the decoder input $z\sim p_{\theta}(z|x_{1,T})$. We then sample with an auto-regressive decoder $p_{\theta}(x^{\prime}_{t+1}|x^{\prime}_{2:t},x_{1,T},z)$. The training process can be found in App. -A.
@@ II-B PerceiverIO encoder
We implement the encoder $q_{\phi}(\operatorname{z}|\operatorname{x}_{1:T})$ via a PerceiverIO model [6], that is, as multi-layered, attention-based neural architecture where each layer is composed of a cross-attention (CA) and as self-attention (SA) block as
$h^{(l+1)}_{1:K}$ $=\mathrm{RMLP}_{\theta}\circ\mathrm{RSA}_{\theta}(h_{1:K}^{(l)}),\quad l=1,\ldots,L$
$h^{(1)}_{1:K}$ $=\mathrm{RMLP}_{\theta}\circ\mathrm{RCA}_{\theta}(h^{(0)}_{1:K},\tilde{x}_{1:T}).$
The initial layer $h^{(0)}_{\theta}$ is a learned parameter, a part of $\theta$. The values $\tilde{x}_{t}=W_{\theta}x_{t}+\tau_{t}$ are embedding the data into the hidden space via a linear layer $W_{\theta}$ and a positional embedding $\tau_{t}$. RCA and RSA denote a residual (R) cross-attention and residual self-attention block defined as $\mathrm{RCA}(h,z)=\mathrm{LayerNorm}(h+\mathrm{CA}(h,z))$ and $\mathrm{RSA}(h)=\mathrm{LayerNorm}(h+\mathrm{SA}(h))$, respectively. The encoder is then defined as a diagonal multivariate Gaussian distribution $q_{\phi}(z|x_{1:T})=\mathcal{N}(z;\mu_{\theta}^{\mathrm{enc}}(h^{(L)}),\sigma_{\theta}^{\mathrm{enc}}(h^{(L)})^{2})$ with $\mu_{\theta}^{\mathrm{enc}}$ and $\sigma_{\theta}^{\mathrm{enc}}$ being simple MLPs. The latent variable $z$ is thus designed to encode the trajectory $x_{1:T},x_{t}\in\mathbb{R}^{d}$ to a shorter sequence of $K$ variables $z_{k}\in\mathbb{R}^{m}$. Note that the choice $K=1$ results in a single vector embedding of the trajectory $x_{1:T}$.
@@
This type of Perceiver architecture is designed to reduce the quadratic ($T\times T$) computational complexity of a Transformer encoder model (self-attention layers) on $x_{1:T}$ by only cross-attending the sequence ($T\times K$) and performing less computationally demanding self-attentions ($K\times K$) on $z_{1:K}$. Generally, there is a large class of Perceiver models/blocks with learnable queries or latents that are used for transforming data between different modalities, for example, a Perceiver model with $K$ queries can be used as a head-block in a classification model with $K$ classes.
@@ II-C Autoregressive decoder
$h^{(l)}_{1:T}=\mathrm{RMLP}^{(l+1)}_{\theta}\circ\mathrm{RCA}_{\theta}^{(l+1)}(\cdot,z)\circ\mathrm{RCSA}_{\theta}^{(l+1)}(h^{(l)}_{1:T})$
with $h^{(0)}_{1:T}=\tilde{x}_{1:T}$ being the embeddings defined in the previous section. The value $h^{(L)}_{t}$ is then used to define the autoregressive model $p(x_{t+1}|x_{2:t},x_{1,T},z)=\mathcal{N}(x_{t+1};\mu_{\theta}^{\mathrm{dec}}(h^{(L)}_{t}),\sigma_{\theta}^{\mathrm{dec}}(h^{(L)}_{t})^{2})$. The additional “C” in the layer definition in $\mathrm{RCSA}$ denotes a casual attention model implemented via triangular attention masks. Unlike Transformers for text infilling [8, 9], we do not include segmentation embedding since the length and position of the start and end points in the model are already provided.
II-D Trajectory generator—Guided decoding
In addition to learning a conditional autoencoder model for trajectory generation we would also like to extend the model to be able to perform online trajectory adaptation, for example, add via-points and obstacle avoidance. To achieve this we do online (approximate) probabilistic inference using $p_{\theta}(x_{t+1}|x_{1:t},x_{T},z)$ without updating the model parameter $\theta$. The obvious choice is a Bayesian approach where the adaptation can be formulated as a Bayesian update of $p_{\theta}(x_{t+1}|x_{1:t},x_{T},z)$, via a likelihood $p(y_{t+1}|x_{t+1})$. Alternatively, the likelihood can also be used to implement constraints, for example, a Bayesian update with a likelihood function $\mathrm{I}_{[a,b]}(x)$ can implement trajectory boundaries.
For generic constraints we can formulate online trajectory adaptation as
@@ II-D Trajectory generator—Guided decoding
where $H(\cdot)$ is the Heaviside function, $c(\cdot)$ is a constraint function. The constraint in (2) expresses what proportion $0\leq\alpha\leq 1$ of the probability mass is allowed not to satisfy the constraint [10]. The KL objective is convex in $q$, the constraints are linear, and the optimal solution $(q^{\ast},\lambda^{\ast})$ satisfies
@@
Solving this optimization, however, even for as Gaussian $q$ and a $c$ implementing a boundary constraint can be numerically demanding. For more general constraints such as obstacle avoidance, (2) can be analytically intractable. For this reason, in the following we will only exploit the form of $q^{\ast}(x_{t+1})$ to propose fast online trajectory adaptation methods. Various trajectory adaptations can be conveniently formulated either as (1) by choosing the appropriate constraint function $c(\cdot)$ or as a Bayesian update with the appropriate likelihood. These will be detailed in the following.
@@ II-E1 Obstacle avoidance with beam search
Beam search (BS) [11] is a classic heuristic in the field of natural language processing (autoregressive models). It is used to generate trajectories with high likelihood or other specific score values. The method recursively generates trajectories as follows. A set of $S$ trajectories $\{x_{2:t}^{(s)}\}$ are stored together with their scores, for example, the path likelihood $\log p(x_{2:t}^{(s)}|x_{1,T},z)$. For each trajectory $S$ new samples $x_{t+1}^{(s,r)}\sim p(x_{t+1}|x_{2:t}^{(s)},x_{1,T},z),r=1,\ldots,S$ are generated as candidates and the corresponding total scores are computed for the extended trajectories $(x^{(s)}_{2:t},x_{t+1}^{(s,r)})$ for all $(s,r)$ pairs. The overall top $S$ extended trajectories are selected for the next step. Inspired by the form of $q^{\ast}(x_{t+1})$, we define the score by adding $\tilde{s}(x_{t+1})=-\gamma\max(c(x_{t+1}),0)$ with $c(x)=\delta^{2}-\|\,x-x_{\mathrm{obstacle}}\,\|^{2}$ to the likelihood.
@@ II-E3 Via points through Brownian bridge
II-E4 Constrained Beam Search for Brownian bridge
@@
We generate samples of $x_{t+1}$ for each sample $x_{t}$, using the Transformer decoder incorporating via point method, i.e., Brownian Bridge. We categorize the generated trajectories into three banks: Bank0 for trajectories that do not meet the constraint; Bank1 for trajectories that have applied the controller but still do not satisfy the constraint; and Bank2 for trajectories that satisfy the constraint. For the trajectory $x_{1:t}$ that satisfies the via-point constraint, we generate a single sample of $x_{t+1}$ using the Transformer decoder exclusively for Bank2. If $x_{1:t}$ fails to satisfy the constraint, we then generate one $x_{t+1}$ sample from the Transformer decoder with the controller. In this case, if $x_{t+1}$ satisfies the constraint, the sample is allocated to Bank2; otherwise, it goes to Bank1. Lastly, in cases where $x_{1:t}$ does not satisfy the constraint, we generate $N_{s}$ $x_{t+1}$ samples from the Transformer decoder only. Here, if $x_{t+1}$ satisfies, the sample is classified into Bank2; if not, into Bank0. Following this, we select the samples from each bank with the order of Bank2, Bank1, Bank0, iteratively based on the negative log-likelihood criterion until we have $N_{s}$ samples in total for $x_{t+1}$.

### v2

# IEEE-RAS International Conference on Humanoid Robots, 2024

@@ Abstract
We present a novel motion generation approach for robot arms, with high degrees of freedom, in complex settings that can adapt online to obstacles or new via points. Learning from Demonstration facilitates rapid adaptation to new tasks and optimizes the utilization of accumulated expertise by allowing robots to learn and generalize from demonstrated trajectories. We train a transformer architecture, based on conditional variational autoencoder, on a large dataset of simulated trajectories used as demonstrations. Our architecture learns essential motion generation skills from these demonstrations and is able to adapt them to meet auxiliary tasks. Additionally, our approach implements auto-regressive motion generation to enable real-time adaptations, as, for example, introducing or changing via-points, and velocity and acceleration constraints. Using beam search, we present a method for further adaption of our motion generator to avoid obstacles. We show that our model successfully generates motion from different initial and target points and that is capable of generating trajectories that navigate complex tasks across different robotic platforms.
@@ I Introduction
Motion generation is a challenging topic for robot arms with high degrees of freedom (DoF). One main challenge is to generate smooth trajectories to solve a given task while ensuring the robot and its environment’s safety, and respecting the robot’s embodiment constraints. This problem is especially complex when the task involves auxiliary objectives beyond reaching a target point, such as avoiding obstacles, passing via specified points, or imposing velocity or acceleration constraints. Another particularly intriguing aspect of motion generation is the incorporation of learning from demonstration (LfD) within the same process. Allowing users or agents to present a guiding trajectory based on prior knowledge helps the rapid development of robotic applications. This capability is crucial for adapting to novel tasks and leveraging accumulated expertise effectively. Motion primitives offer a promising avenue for achieving this goal [1, 2, 3, 4]. By encapsulating fundamental motion behaviors into primitives, the robot can learn and generalize from demonstrations, promoting versatility and adaptability in complex tasks. Integrating learning from demonstration with the challenges of motion generation presents an exciting direction for advancing the capabilities of robot arms with high degrees of freedom.
Traditional approaches to motion primitives often rely on a limited number of demonstrations, focusing primarily on adapting learned motions [2, 3]. However, these methods struggle with generalization when applied to new or variations of the learned tasks that go beyond simple interpolation. Our work seeks to address these limitations by leveraging generative models, specifically an autoregressive framework, which allows learning from extensive datasets and significantly improves generalization. Unlike diffusion-based models, our autoregressive model facilitates online motion adaptation. Diffusion-based models [5] require multiple iterations to generate a full trajectory and, therefore, can not adapt fast to changes in the environment or the robot embodiment. In contrast our autoregressive model supports the integration of online feedback, allowing it to adapt to changes to the task, e.g. different via-points, obstacle locations, or the robot, e.g. acceleration constraints arises from interacting with physical objects.
@@
In this work, we propose a framework for learning motion generation. We create a large dataset of robot motion trajectories with multiple simulated robots and train a transformer architecture to learn to generate motion based on given initial and target points. We design a Transformer conditional variational autoencoder. The trained generative model encapsulates basic motion generation skills based on the training data. We further propose an approach to efficiently adapt the generated motion to satisfy auxiliary tasks and constraints such as obstacle avoidance, via points and velocity constraints. The process is auto-regressive allowing for integrating feedback from the real system during motion generation. Our approach allows for generating motion based on an initial and target point only, or to imitate a full trajectory given by a user or external agent. Our experiments show the usability of our method for imitating and adapting user-given trajectories to a given task, such as via point constraints, obstacles, and velocity and acceleration constraints. We also show the benefits of training with data collected using various robots.
@@ II Methods
We design a model to encode and generate robot trajectories by leveraging a transformer-based architecture. We generate trajectories in an autoregressive manner, allowing for real-time adaptation to constraints such as obstacle avoidance and velocity limits. In this section, we present a Transformer-based autoencoder model with an autoregressive decoder and the online adaptation methods that are tailored to this decoder.
@@ II-A Inference model for the infilling task
where the latent variables (representations) $z$ are expected to encode features of paths with starting point $x_{1}$ and endpoint $x_{T}$. Since the likelihood is analytically intractable, we learn this model by approximate maximum likelihood via a variational bound. We minimize the negative evidence lower bound (ELBO)
@@
Here $p_{\theta}(z|x_{1,T})$ is the conditional prior model, $p_{\theta}(x_{2:T-2}|x_{1,T},z)$ is the decoder/generation model, and $q_{\phi}(z|\operatorname{x}_{1:T})$ is the encoder/recognition or (variational) approximate posterior distribution model. The training method can be found in App. -A. To generate a new trajectory, we sample a latent feature $z$ from the learnt prior $z\sim p_{\theta}(z|x_{1,T})$. We then use the feature $z$ as a conditioning variable in the learnt decoder model $p_{\theta}(x^{\prime}_{2:T-2}|x_{1,T},z)$.
In this paper we use an autoregressive decoder model $p_{\theta}(x^{\prime}_{2:T-2}|x_{1,T},z)=\prod_{t}p_{\theta}(x^{\prime}_{t+1}|x^{\prime}_{2:t},x_{1,T},z)$. We do this because it is a more natural choice for doing online trajectory guidance and adaptation compared to methods that generate a whole trajectory like, for example, diffusion probabilistic models [7].
@@ II-B PerceiverIO encoder
We define an encoder $q_{\phi}(\operatorname{z}|\operatorname{x}_{1:T})$ that is architecturally similar to a PerceiverIO model [8], that is, a multi-layered, attention-based neural architecture where the initial block uses cross-attention (CA) layer and the subsequent blocks use a shared self-attention (SA) layer
$h^{(1)}_{1:K}$ $=\mathrm{RMLP}_{\phi}\circ\mathrm{RCA}_{\phi}(h^{(0)}_{1:K},\tilde{x}_{1:T})$
$h^{(l+1)}_{1:K}$ $=\mathrm{RMLP}_{\phi}\circ\mathrm{RSA}_{\phi}(h_{1:K}^{(l)}),\quad l=1,\ldots,L$
The initial layer $h^{(0)}_{\phi}$ is a learned parameter, a part of $\phi$. The values $\tilde{x}_{t}=Wx_{t}+\tau_{t}$ are embedding the data into the hidden space via a linear layer $W$ and a positional embedding $\tau_{t}$. RCA and RSA denote a residual (R) cross-attention and a residual self-attention block defined as $\mathrm{RCA}(h,z)=\mathrm{LayerNorm}(h+\mathrm{CA}(h,z))$ and $\mathrm{RSA}(h)=\mathrm{LayerNorm}(h+\mathrm{SA}(h))$, respectively. RMLP denotes a residual MLP block with one hidden layer. Note that the resulting self-attention block is a shared block. The encoder is then defined as a diagonal multivariate Gaussian distribution $q_{\phi}(z|x_{1:T})=\mathcal{N}(z;\mu_{\phi}^{\mathrm{enc}}(h^{(L)}),\sigma_{\phi}^{\mathrm{enc}}(h^{(L)})^{2})$ with $\mu_{\phi}^{\mathrm{enc}}$ and $\sigma_{\phi}^{\mathrm{enc}}$ being simple MLPs. The latent variable $z$ is thus designed to encode the trajectory $x_{1:T},x_{t}\in\mathbb{R}^{d}$ to a shorter sequence of $K$ variables $z=\{z_{k}\}_{k=1}^{K},z_{k}\in\mathbb{R}^{m}$ ($K$ is a chosen hyper-parameter). Note that the choice $K=1$ results in a single vector embedding of the trajectory $x_{1:T}$.
@@
This type of Perceiver architecture is designed to reduce the quadratic ($T\times T$) computational complexity of a Transformer encoder model (self-attention layers) on $x_{1:T}$ by cross-attending the sequence ($T\times K$) and performing less computationally demanding self-attentions ($K\times K$) on $z_{1:K}$. Generally, there is a large class of Perceiver models/blocks with learnable queries or latents that are used for transforming data between different modalities/embeddig sizes, for example, a Perceiver model with $K$ queries can be used as a head-block in a classification model with $K$ classes.
@@ II-C Autoregressive decoder
$h^{(l+1)}_{1:T}=\mathrm{RMLP}^{(l+1)}_{\theta}\circ\mathrm{RCA}_{\theta}^{(l+1)}(\cdot,z)\circ\mathrm{RCSA}_{\theta}^{(l+1)}(h^{(l)}_{1:T})$
with $h^{(0)}_{1:T}=\tilde{x}_{1:T}$ being the embeddings as defined in the previous section. The data $x_{1,T}$ together with the representation $z$ are direct inputs to the decoder with which the trajectory generation starts. To allow for flexible output length, we do not fix the position of $x_{T}$ at the end; instead, we set it as the first step of the input. The value $h^{(L)}_{t}$ is then used to define the autoregressive model $p(x_{t+1}|x_{2:t},x_{1,T},z)=\mathcal{N}(x_{t+1};\mu_{\theta}^{\mathrm{dec}}(h^{(L)}_{t}),\sigma_{\theta}^{\mathrm{dec}}(h^{(L)}_{t})^{2})$. The additional “C” in the layer definition in $\mathrm{RCSA}$ denotes a casual attention model implemented via triangular attention masks. Unlike in Transformers models for text infilling [10, 11], we do not need to include segmentation embedding since the length and position of the start and end points in the model are already provided.
II-D Training-free trajectory generator—adaptive decoding
In addition to learning a conditional autoencoder model for trajectory generation, we would also like to extend the model to be able to perform online trajectory adaptation, for example, to add via-points and obstacle avoidance. To achieve this, we do online (approximate) probabilistic inference using $p_{\theta}(x_{t+1}|x_{1:t},x_{T},z)$ without updating the model parameter $\theta$. The obvious choice is a Bayesian approach where the adaptation can be formulated as a Bayesian update of $p_{\theta}(x_{t+1}|x_{1:t},x_{T},z)$, via a likelihood $p(y_{t+1}|x_{t+1})$. Alternatively, the likelihood can also be used to implement constraints, for example, a Bayesian update with a likelihood function $\mathrm{I}_{[a,b]}(x)$ can implement trajectory boundaries $a_{i}\leq x_{i}\leq b$.
To use generic constraints, we can formulate online trajectory adaptation as the constrained optimisation problem
@@ II-D Training-free trajectory generator—adaptive decoding
where $H(\cdot)$ is the Heaviside function and $c(\cdot)$ is a desired constraint function. The constraint in (2) expresses what proportion $0\leq\alpha\leq 1$ of the probability mass is allowed not to satisfy the constraint, see for example [12]. The KL objective is convex in $q$, the constraints are linear, and the optimal solution $(q^{\ast},\lambda^{\ast})$ satisfies
@@
Solving this optimization, however, even for a Gaussian $q$ and a constraint $c$ implementing simple boundary constraints can be numerically demanding. For more general constraints such as obstacle avoidance, (2) can be analytically intractable. For this reason, in the following we will only exploit the form of $q^{\ast}(x_{t+1})$ to propose fast online trajectory adaptation methods. Various trajectory adaptations can be conveniently formulated either as (1) by choosing the appropriate constraint function $c(\cdot)$ or as a Bayesian update with the appropriate likelihood. These will be detailed in the following.
@@ II-E1 Obstacle avoidance with beam search
Beam search (BS) [13] is a classic heuristic in the field of natural language processing (autoregressive models). It is used to generate trajectories with high likelihood or other specific score values. The method recursively generates trajectories as follows. A set of $S$ trajectories $\{x_{2:t}^{(s)}\}$ are stored together with their scores, for example, the path likelihood $\log p(x_{2:t}^{(s)}|x_{1,T},z)$. For each trajectory $S$ new samples $x_{t+1}^{(s,r)}\sim p(x_{t+1}|x_{2:t}^{(s)},x_{1,T},z),r=1,\ldots,S$ are generated as candidates and the corresponding total scores are computed for the extended trajectories $(x^{(s)}_{2:t},x_{t+1}^{(s,r)})$ for all pairs $(s,r)$. The overall top $S$ extended trajectories are selected for the next step. Inspired by the form of $q^{\ast}(x_{t+1})$, we define the score by adding $\tilde{s}(x_{t+1})=-\gamma\max(c(x_{t+1}),0)$ with $c(x)=\delta^{2}-\|\,x-x_{\mathrm{obstacle}}\,\|^{2}$ to the likelihood.
@@ II-E3 Via points through Brownian bridge
II-E4 Constrained Beam Search for Brownian bridge (optional)
@@
We generate samples of $x_{t+1}$ for each sample $x_{t}$, using the Transformer decoder incorporating via point method, i.e., Brownian Bridge. We categorize the generated trajectories into three banks: Bank-0 for trajectories that do not meet the constraint; Bank-1 for trajectories that have applied the controller but still do not satisfy the constraint; and Bank-2 for trajectories that satisfy the constraint. For the trajectory $x_{1:t}$ that satisfies the via-point constraint, we generate a single sample of $x_{t+1}$ using the Transformer decoder exclusively for Bank-2. If $x_{1:t}$ fails to satisfy the constraint, we then generate one $x_{t+1}$ sample from the Transformer decoder with the controller. In this case, if $x_{t+1}$ satisfies the constraint, the sample is allocated to Bank-2; otherwise, it goes to Bank-1. Lastly, in cases where $x_{1:t}$ does not satisfy the constraint, we generate $N_{s}$ $x_{t+1}$ samples from the Transformer decoder only. Here, if $x_{t+1}$ satisfies, the sample is classified into Bank-2; if not, into Bank-0. Following this, we select the samples iteratively from each bank with the order of Bank-2, Bank-1, Bank-0, based on the negative log-likelihood criterion until we have $N_{s}$ samples in total for $x_{t+1}$.
@@ III Related work
Recently, diffusion models [7] for motion generation [5] have been developed. However, these models generally require a large number of iterations to transform Gaussian noise into usable data, which is computationally intensive. In addition, it can only produce complete sequence data. In contrast, our model does not require multiple iterations for completion and is an autoregressive model that generates data for each time step which is more useful for online motion adaption.